meena is thice as old as sheraz. fatima will be twice as old as meena 6 years hence . six years ago sheraz was five years old . what is fatimas present age.
1
Expert's answer
2012-11-12T12:32:04-0500
Let's denote the Meena's, Fatima's and Sheraz's ages by M, F and S respectfully. Then we've got the following system of equations:

M = 3S& (1)
F+6 = 2M (2)
S-6 = 5& (3)

(3) ==> S = 5+6 = 11;
(1): M = 3S = 3*11 = 33;
(2) ==> F = 2M - 6 = 2*33 - 6 = 66-6 = 60.

So, Fatima's present age is 60 years.